OpenStudy (nickymail48):

8/ (-0.02) easy way to know how to solve

OpenStudy (jdoe0001):

have you covered the exponents rules yet?

OpenStudy (nickymail48):

need some review please

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg <--- any ideas on how to convert the numerator to an exponential form?

OpenStudy (nickymail48):

yah forgot all those rules

OpenStudy (jdoe0001):

hmmm what are you supposed to do with the equation anyway?

OpenStudy (nickymail48):

its asking to divide

OpenStudy (jdoe0001):

\(\bf \cfrac{8}{-0.02}\quad \begin{cases} 8\to 2\cdot 2\cdot 2\\ 0.\underline{02}\to\frac{0\underline{02}}{1\underline{00}}\to \frac{2}{100}\to 2\cdot \frac{1}{100}\to 2\cdot 10^{-2} \end{cases}\qquad thus \\ \quad \\ \cfrac{8}{-0.02}\implies \cfrac{\cancel{ 2 }\cdot 2\cdot 2}{-\cancel{ 2 }\cdot 10^{-2}}\)

OpenStudy (nickymail48):

so its -400? i dont get it

OpenStudy (jdoe0001):

well... notice....all I did was factor the top and bottom the bottom by making it a 10base type so \(\large { \cfrac{8}{-0.02}\quad \begin{cases} 8\to 2\cdot 2\cdot 2\\ 0.\underline{02}\to\frac{0\underline{02}}{1\underline{00}}\to \frac{2}{100}\to 2\cdot \frac{1}{100}\to 2\cdot 10^{-2} \end{cases} \\ \quad \\ \cfrac{8}{-0.02}\implies \cfrac{\cancel{ 2 }\cdot 2\cdot 2}{-\cancel{ 2 }\cdot 10^{-2}}\implies \cfrac{4}{-10^{-2}}\implies \cfrac{4}{-0.01} }\)

OpenStudy (nickymail48):

right its looking for a whole number but i dont know how to read the question

OpenStudy (jdoe0001):

a whole number? well... you can just do a division of that... that'd give you actually -400

OpenStudy (jdoe0001):

unless you mean to do it by long division

OpenStudy (nickymail48):

yes! but where do i start with dividing?

OpenStudy (jdoe0001):

say 8/-0.02 so..... you'd notice first how many decimals on either the dividend and divisor and add that many zeros to the other, while dropping the decimal point say in this case dividend 8 has no decimals.... so no zeros added to the divisor divisor -0.02 has 2 decimals so we add 2 zeros to 8 so it ends up looking like \(\large \cfrac{8}{-0.02}\implies 8\div -0.\underline{02}\implies 8\underline{00}\div -\quad 2\)

OpenStudy (jdoe0001):

once we dropped the decimal point from -0.02 we would end up with -002 which is in significant numbers only -2